# Lil: Less is Less When Applying Post-Training Sparse-Attention Algorithms in Long-Decode Stage

### v1

@@ 1 Introduction
Large language models (LLMs) openai2024o1; damai2024deepseek exhibit strong capabilities across a wide range of complex tasks, such as automatic code generation wang2023how and creative writing yushi2024longbench. Users interact with LLMs through natural-language-based prompts—sequences of tokens [^1] Tokens may be thought of roughly as words.. This combination of usability and capability has driven rapid and widespread adoption. As a result, LLMs must now be deployed at scale to handle an increasingly large and diverse set of requests, including long-input requests (e.g., document-question answering bai2024longbench; yushi2024longbench; longformer; wu2025loongserve), long-output requests (e.g., chain-of-thought reasoning aime; math500; yao2023react, long-form writing yushi2024longbench, code generation wang2023how), as well as requests requiring both wu2025resum. Longer inputs and outputs significantly increase inference latency and resource consumption, posing substantial challenges for large-scale deployment.
To address these inference challenges, prior work typically decomposes inference into two stages: prefill and decode. In the prefill stage, the model processes tokens given by users. It computes the Key (K) and Value (V) vectors for all tokens, stores these vectors in the KV cache, and generates the first output token to initiate the decode stage. In the decode stage, the model iteratively processes each newly generated token. It computes the KV vectors for the new token, appends these vectors to the KV cache, and generates the next token. This process repeats until a specified stopping criterion is met. This paper focuses on accelerating the decode stage, which dominates total inference time hu2025raas; yao2023react, especially in reasoning-intensive tasks.
To optimize the decode stage, a major line of work introduces sparse-attention algorithms [^2] Although sparse attention also applies to prefill, this work focuses on the decode stage. We also emphasize post-training sparsity. Training-aware sparsity algorithms such as DeepSeek NSA (Native Sparse Architecture) nsa2025yuan fall outside our scope and are discussed in related work., aiming to reduce both time and memory complexity zhang2023h2o; xiao2023sink; tang2024quest; hu2025raas; chen2024arkvale. First, sparse attention reduces time complexity. Full attention requires each decode token to attend to all previous tokens. In contrast, sparse attention requires each decode token to attend to only the top-k most relevant tokens, substantially reducing computation while maintaining accuracy. Second, sparse attention may reduce memory complexity. Some algorithms reduce memory by discarding irrelevant KV vectors during decoding zhang2023h2o; xiao2023sink; hu2025raas, while others retain the full KV cache tang2024quest; chen2024arkvale, but therefore cannot reduce the memory footprint.
@@
We implement a unified framework that supports the integration and comparative evaluation of diverse sparse-attention algorithms. We also incorporate Guardian into this framework. Our evaluation yields two key findings (Section 6). First, on reasoning-intensive benchmarks math500; aime; gsm8k, Guardian reduces total token wastage by up to 90% compared to decoding without early stopping, with less than 2% accuracy drop. Second, experiments show that Guardian can also be applied to general cases of prolonged Chain-of-Thought (CoT) generation even without sparse-attention algorithms [^4] Prolonged CoT arises from ill reasoning patterns induced by data quality issues, human preference biases for long sentences, or reward hacking, rather than from the information-loss-and-reconstruction characteristic of the Lil problem; despite related, it lies beyond the scope of this work..
@@ 2.2 Cost of Long-Prefill and Long-Decode Inference
Both long-prefill and long-decode inference incur substantial memory and time overheads. In terms of memory, processing 128k tokens with the LLaMA 3.1 8B model in FP16 requires up to 16 GB, in addition to the 16 GB of model parameters [^5] https://huggingface.co/blog/llama31. In terms of time, inference on 32k tokens can take from tens to thousands of seconds on vLLM 0.6.1 with the same model hu2025raas, with large variance driven primarily by the decode stage: longer generations incur proportionally higher latency.
The decode stage dominates end-to-end inference time, especially for reasoning-intensive tasks openai2024o1; wang2024openr; zhao2024marco; wei2022chain. For instance, the OpenAI o1 openai2024o1 may spend tens to hundreds of seconds in internal “thinking” before producing a final answer [^6] https://www.reddit.com/r/OpenAI/comments/1frdwqk/your_longest_thinking_time_gpt4_o1_o1mini/. Accordingly, this paper focuses on optimizing long-decode inference.
@@ 2.3 Post-Training Sparse-Attention Algorithms for Long-Decode Optimization
To reduce memory and time complexity in LLM inference, a substantial body of work proposes sparse-attention algorithms xiao2023sink; zhang2023h2o; tang2024quest; chen2024arkvale; nsa2025yuan, which restrict attention computation at each decode step to a small subset of critical tokens—often fewer than 10% of the context tang2024quest. First, sparse attention reduces time complexity. Full attention requires each decode token to attend to all previous tokens. In contrast, sparse attention requires each decode token to attend to only the top-k most relevant tokens. Second, sparse attention may reduce memory complexity. Some algorithms reduce memory by discarding irrelevant KV vectors during decoding, such as H2O and Sink zhang2023h2o; xiao2023sink; hu2025raas, while others retain the full KV cache, such as infLLM and Quest tang2024quest; chen2024arkvale; infllm2024xiao, but therefore cannot reduce the memory footprint.
Sparse-attention algorithms can be broadly categorized into two types. Training-aware algorithms incorporate sparsity directly into the model architecture and training procedure. Despite being effective, they require architectural modifications and incur substantial training costs, and their sparsity is difficult to disable once deployed. Representative examples include DeepSeek Native Sparse Attention (NSA) and DeepSeek Sparse Attention (DSA) damai2024deepseek; liu2024deepseek-v3. In contrast, post-training algorithms apply sparsity to fully trained dense models at inference time. These algorithms are plug-and-play and training-free, and reportedly have demonstrated strong performance in preserving accuracy while reducing latency and memory consumption. This paper focuses on Post-Training Sparse-attention algorithms in the Decode stage (PTSD).
@@ 4 Compression Theory
To this end, we adopt LZ77 ziv1977lz77, a compression algorithm, which is simple, efficient, and grounded in theory. First, the key insight of LZ77 is simple: replace repeated substrings with concise references (e.g., offset-length pairs) to their earlier occurrences. When later segments of a sequence frequently recur in earlier segments, the resulting compression ratio (compressed length divided by original length) is small. Second, LZ77 is computationally efficient: compressing a sequence of 128k tokens takes approximately 34 ms (Section 6), which is comparable to the time to decode a single token zheng2024sglang. Third, LZ77 admits strong theoretical guarantees. The achieved compression ratio $\rho$ satisfies
@@ 6.1 Experimental Setup
Implementation. We implement Guardian based on Hugging Face hf with 2k lines of Python code. We port the LZ77 algorithm from the famous compression tool Gzip [^7] https://www.gzip.org/.
Datasets. We take the first 200 test cases from each of the following three open-source datasets for our benchmarks: GMS8K gsm8k, MATH500 math500, and AIME aime, to test the reasoning ability of language models. First, GMS8k gsm8k contains 8.5k high-quality, linguistically diverse grade-school math problems. These human-written problems need solutions that involve multi-step reasoning and a series of basic arithmetic operations. Second, MATH500 math500 contains 500 challenging problems sourced from high school math competitions with five distinct levels based on the Art of Problem Solving (AoPS) framework, ranging from level 1 to level 5. Third, AIME aime is a math problem dataset collected from the American Invitational Mathematics Examination (AIME) competition from 1983 to 2024, designed to challenge the most exceptional high school math students in the United States. These problems cover various fields, such as algebra, geometry, and number theory.
Models. We evaluate our algorithm using three popular models: DeepScaleR-1.5B-Preview [^8] https://pretty-radio-b75.notion.site/DeepScaleR-Surpassing-O1-Preview-with-a-1-5B-Model-by-Scaling-RL-19681902c1468005bed8ca303013a4e2, DeepSeek-R1-Distill-Llama-8B deepseekai2025deepseekr1, and Qwen1.5-MoE-A2.7B-Chat [^9] https://qwenlm.github.io/blog/qwen-moe/. These models span diverse architectures (dense vs. MoE), training recipes, and parameter scales. In this paper, we use DSR to denote DeepScaleR-1.5B-Preview, DSL to denote DeepSeek-R1-Distill-Llama-8B, and Qwe to denote Qwen1.5-MoE-A2.7B-Chat.
Metrics. We evaluate efficiency and accuracy using two metrics. Token savings is defined as the ratio of the number of tokens generated without Guardian to that generated after applying Guardian. Accuracy wang2024openr measures the mathematical equivalence between an LLM’s output and the ground-truth answer. For each test case, it is either correct or incorrect, and the overall accuracy is reported as the percentage of correctly solved problems across the entire dataset. When reporting accuracy changes, we use the absolute difference in accuracy, expressed in percentage points.
@@ 6.2 Evaluation Results
Cost of LZ77. We generate random strings of up to 128k tokens and measure the cost of LZ77 compression, which is approximately 34 ms. Because random strings exhibit minimal repetition and are more expensive to compress than natural language text, and because 128k tokens far exceed the sequence lengths used in our experiments, this measurement represents an upper bound. The resulting cost is comparable to the latency of decoding a single token zheng2024sglang. With $f=250$, compression is invoked once every 250 decoding steps, making the overall overhead negligible.
@@ 7.1 Sparse-Attention Algorithms
Sparse-attention algorithms exploit the fact that only a small subset of tokens receives significant attention scores, and we categorize them along two dimensions. (1) Inference stage. Some algorithms primarily accelerate the prefill stage, such as MInference jiang2024minference, FlexPrefill lai2025flexprefill, and SpargeAttention zhang2025spargeattn, while others focus on the decode stage, including H2zhang2023h2o, StreamingLLM xiao2023sink, Quest tang2024quest, InfLLM infllm2024xiao, DuoAttention xiao2024duoattention, and RaaS hu2025raas. (2) Training dependency. Training-aware algorithms incorporate sparsity into the model architecture and are trained jointly with the model, such as DeepSeek NSA nsa2025yuan and DSA liu2024deepseek-v3. In contrast, training-free algorithms operate as plug-ins and can be readily applied to pretrained models; examples include H2O, StreamingLLM, Quest, InfLLM, DuoAttention, and RaaS.
@@ 7.2 Chain-of-Thoughts Compression
Prior work on CoT compression, often referred to as long-to-short reasoning, can be broadly categorized into two classes. Training-aware approaches integrate compression into the training process, including ThinkPrune hou2025thinkprune, DLER liu2025dler, and O1-Pruner luo2025o1pruner. In contrast, post-training approaches operate on pretrained models, such as Answer5 liu2025answer5, HALT-COT sun2025haltcot, and UnCertT zhu2025uncert.
@@ 8 Conclusion
Evaluation on a limited set of datasets and models. Our evaluation covers only three models and three datasets. As such, the results may not generalize beyond these specific configurations. Although models with longer context lengths (e.g., Qwen2.5-Max, DeepSeek-r1) and datasets such as GPQA Diamond and Codeforces exist, exhaustive evaluation across all combinations is computationally prohibitive hu2023pcrml. As reported in prior work zhong2024distserve, decoding a single token can take approximately 30 ms; thus, processing 16k tokens on an A100-80GB GPU requires around 8 minutes. Running 200 test cases would take over a day on a single GPU, making large-scale evaluation infeasible with limited resources. Despite these constraints, we select datasets spanning three levels of difficulty and models covering diverse architectures (dense vs. MoE), training recipes, and parameter scales. We therefore believe that the Lil problem is not an artifact of specific configurations but is universal across datasets and models.
Evaluation on a limited set of sparse-attention algorithms. Our evaluation covers only four sparse-attention algorithms, and thus the results may not directly generalize beyond this set. Nevertheless, we contend that the Lil problem is inherent to sparse-attention algorithms, as all assume sparsity patterns—i.e., that a few tokens are important—and may lose information (Section 4). We select diverse algorithms, including the pioneering H2O, the widely-used streamingLLM (recently applied in GPT-OSS openai2025gptoss), the state-of-the-art Quest, and its counterpart infLLM. Other algorithms, such as clusterKV and PQCache, differ only in how K vectors are grouped and do not alter the use of sparsity. Hence, we believe the Lil problem is not specific to configurations but universal across sparse-attention algorithms.
Figure 4: Illustration of the LZ77 algorithm.
Figure 5: Examples of information loss and gain.
GSM8K MATH-500 AIME
% H2O Sink Quest infLLM H2O Sink Quest infLLM H2O Sink Quest infLLM
DSR 128 83.0/67.7 59.3/68.3 0.9/7.7 3.8/6.8 77.4/82.0 79.7/84.3 7.3/11.7 3.0/8.8 89.1/93.7 84.8/92.6 37.2/18.2 17.6/14.1
256 55.9/62.3 27.0/68.4 1.5/18.2 1.1/7.1 68.9/78.7 72.5/80.2 3.4/13.7 1.2/11.1 92.3/91.2 0.0/90.9 0.0/20.0 28.0/16.6
512 23.7/57.9 20.1/64.9 0.3/25.7 0.0/13.2 52.0/67.5 32.0/73.6 2.1/18.4 3.1/14.3 78.1/86.5 90.6/87.8 1.2/24.5 5.2/25.0
1024 15.1/61.9 6.8/63.9 0.0/50.4 0.5/41.4 26.6/64.8 16.7/65.7 0.9/23.6 1.0/26.2 49.8/79.9 30.7/80.6 0.6/25.0 5.0/36.9
DSL 128 21.6/22.8 3.8/8.6 0.0/2.6 0.0/0.5 83.9/75.2 41.5/48.3 19.8/14.3 1.9/4.0 91.9/91.0 59.2/71.0 30.7/23.1 0.0/4.9
256 17.8/23.8 0.6/19.5 0.3/7.8 0.0/1.4 68.8/76.1 16.6/51.5 12.4/26.3 0.0/2.4 96.9/87.8 88.3/72.4 21.2/45.7 0.0/4.1
512 4.9/29.5 1.1/16.4 0.7/6.0 0.0/0.9 46.5/74.2 7.7/49.7 10.4/35.0 0.1/6.8 73.5/87.6 22.0/69.3 33.8/55.8 6.7/10.4
1024 3.7/20.5 0.8/10.0 0.4/4.0 0.0/1.0 21.0/71.7 8.2/52.0 7.4/42.5 1.1/12.2 60.3/84.2 23.4/69.3 21.6/54.2 2.3/20.7
Qwe 128 0.0/0.9 0.0/1.2 0.0/0.0 0.0/0.0 9.6/17.0 7.7/14.6 0.0/0.2 0.0/0.0 66.3/39.3 66.4/36.5 0.0/1.4 0.0/0.2
256 0.0/0.8 0.0/2.5 0.0/0.0 0.0/0.0 7.0/21.2 2.0/15.4 0.0/0.0 0.0/0.5 48.6/42.6 0.0/31.7 0.0/1.2 0.0/0.8
512 0.0/2.8 0.0/0.3 0.0/0.0 0.0/0.0 6.5/14.2 0.0/7.5 0.0/0.2 0.0/0.5 0.0/32.0 0.0/16.2 0.0/0.6 0.0/1.1
1024 0.0/0.5 0.0/0.0 0.0/0.0 0.0/0.0 2.0/3.6 0.0/1.0 0.0/0.2 0.9/0.5 0.0/11.2 0.0/6.2 9.0/1.4 0.0/2.7
Table 3: Average token savings for correct and incorrect cases. Notation follows Table 1. Each cell reports (token savings on correct cases) / (token savings on incorrect cases).

### v2

@@ 1 Introduction
Large language models (LLMs) OpenAI ; Dai et al. (2024) exhibit strong capabilities across a wide range of complex tasks, such as automatic code generation Wang et al. (2023) and creative writing Bai et al. (2024b). Users interact with LLMs through natural-language-based prompts—sequences of tokens [^1] Tokens may be thought of roughly as words.. This combination of usability and capability has driven rapid and widespread adoption. As a result, LLMs must now be deployed at scale to handle an increasingly large and diverse set of requests, including long-input requests (e.g., document-question answering Bai et al. (2024a, b); Beltagy et al. (2020); Wu et al. (2024)), long-output requests (e.g., chain-of-thought reasoning AIME ; Hendrycks et al. (2021); Yao et al. (2023), long-form writing Bai et al. (2024b), code generation Wang et al. (2023)), as well as requests requiring both Wu et al. (2025). Longer inputs and outputs significantly increase inference latency and resource consumption, posing substantial challenges for large-scale deployment.
To address these inference challenges, prior work typically decomposes inference into two stages: prefill and decode. In the prefill stage, the model processes tokens given by users. It computes the Key (K) and Value (V) vectors for all tokens, stores these vectors in the KV cache, and generates the first output token to initiate the decode stage. In the decode stage, the model iteratively processes each newly generated token. It computes the KV vectors for the new token, appends these vectors to the KV cache, and generates the next token. This process repeats until a specified stopping criterion is met. This paper focuses on accelerating the decode stage, which dominates total inference time Hu et al. (2024); Yao et al. (2023), especially in reasoning-intensive tasks.
To optimize the decode stage, a major line of work introduces sparse-attention algorithms [^2] Although sparse attention also applies to prefill, this work focuses on the decode stage. We also emphasize post-training sparsity. Training-aware sparsity algorithms such as DeepSeek NSA (Native Sparse Architecture) Yuan et al. (2025) fall outside our scope and are discussed in related work., aiming to reduce both time and memory complexity Zhang et al. (2023); Xiao et al. (2024b); Tang et al. (2024); Hu et al. (2024); Chen et al. (2024). First, sparse attention reduces time complexity. Full attention requires each decode token to attend to all previous tokens. In contrast, sparse attention requires each decode token to attend to only the top-k most relevant tokens, substantially reducing computation while maintaining accuracy. Second, sparse attention may reduce memory complexity. Some algorithms reduce memory by discarding irrelevant KV vectors during decoding Zhang et al. (2023); Xiao et al. (2024b); Hu et al. (2024), while others retain the full KV cache Tang et al. (2024); Chen et al. (2024), but therefore cannot reduce the memory footprint.
@@
We implement a unified framework that supports the integration and comparative evaluation of diverse sparse-attention algorithms. We also incorporate Guardian into this framework. Our evaluation yields two key findings (Section 6). First, on reasoning-intensive benchmarks Hendrycks et al. (2021); AIME ; Cobbe et al. (2021), Guardian reduces total token wastage by up to 90% compared to decoding without early stopping, with less than 2% accuracy drop. Second, experiments show that Guardian can also be applied to general cases of prolonged Chain-of-Thought (CoT) generation even without sparse-attention algorithms [^4] Prolonged CoT arises from ill reasoning patterns induced by data quality issues, human preference biases for long sentences, or reward hacking, rather than from the information-loss-and-reconstruction characteristic of the Lil problem; despite related, it lies beyond the scope of this work..
@@ 2.2 Cost of Long-Prefill and Long-Decode Inference
Both long-prefill and long-decode inference incur substantial memory and time overheads. In terms of memory, processing 128k tokens with the LLaMA 3.1 8B model in FP16 requires up to 16 GB, in addition to the 16 GB of model parameters [^5] https://huggingface.co/blog/llama31. In terms of time, inference on 32k tokens can take from tens to thousands of seconds on vLLM 0.6.1 with the same model Hu et al. (2024), with large variance driven primarily by the decode stage: longer generations incur proportionally higher latency.
The decode stage dominates end-to-end inference time, especially for reasoning-intensive tasks OpenAI ; Wang et al. (2024); Zhao et al. (2024); Wei et al. (2022). For instance, the OpenAI o1 OpenAI may spend tens to hundreds of seconds in internal “thinking” before producing a final answer [^6] https://www.reddit.com/r/OpenAI/comments/1frdwqk/your_longest_thinking_time_gpt4_o1_o1mini/. Accordingly, this paper focuses on optimizing long-decode inference.
@@ 2.3 Post-Training Sparse-Attention Algorithms for Long-Decode Optimization
To reduce memory and time complexity in LLM inference, a substantial body of work proposes sparse-attention algorithms Xiao et al. (2024b); Zhang et al. (2023); Tang et al. (2024); Chen et al. (2024); Yuan et al. (2025), which restrict attention computation at each decode step to a small subset of critical tokens—often fewer than 10% of the context Tang et al. (2024). First, sparse attention reduces time complexity. Full attention requires each decode token to attend to all previous tokens. In contrast, sparse attention requires each decode token to attend to only the top-k most relevant tokens. Second, sparse attention may reduce memory complexity. Some algorithms reduce memory by discarding irrelevant KV vectors during decoding, such as H2O and Sink Zhang et al. (2023); Xiao et al. (2024b); Hu et al. (2024), while others retain the full KV cache, such as infLLM and Quest Tang et al. (2024); Chen et al. (2024); Xiao et al. (2024a), but therefore cannot reduce the memory footprint.
Sparse-attention algorithms can be broadly categorized into two types. Training-aware algorithms incorporate sparsity directly into the model architecture and training procedure. Despite being effective, they require architectural modifications and incur substantial training costs, and their sparsity is difficult to disable once deployed. Representative examples include DeepSeek Native Sparse Attention (NSA) and DeepSeek Sparse Attention (DSA) Dai et al. (2024); Liu et al. (2024). In contrast, post-training algorithms apply sparsity to fully trained dense models at inference time. These algorithms are plug-and-play and training-free, and reportedly have demonstrated strong performance in preserving accuracy while reducing latency and memory consumption. This paper focuses on Post-Training Sparse-attention algorithms in the Decode stage (PTSD).
@@ 4 Compression Theory
To this end, we adopt LZ77 Ziv and Lempel (1977), a compression algorithm, which is simple, efficient, and grounded in theory. First, the key insight of LZ77 is simple: replace repeated substrings with concise references (e.g., offset-length pairs) to their earlier occurrences. When later segments of a sequence frequently recur in earlier segments, the resulting compression ratio (compressed length divided by original length) is small. Second, LZ77 is computationally efficient: compressing a sequence of 128k tokens takes approximately 34 ms (Section 6), which is comparable to the time to decode a single token Zheng et al. (2024). Third, LZ77 admits strong theoretical guarantees. The achieved compression ratio $\rho$ satisfies
@@ 6.1 Experimental Setup
Implementation. We implement Guardian based on Hugging Face Hugging Face with 2k lines of Python code. We port the LZ77 algorithm from the famous compression tool Gzip [^7] https://www.gzip.org/.
Datasets. We take the first 200 test cases from each of the following three open-source datasets for our benchmarks: GMS8K Cobbe et al. (2021), MATH500 Hendrycks et al. (2021), and AIME AIME , to test the reasoning ability of language models. First, GMS8k Cobbe et al. (2021) contains 8.5k high-quality, linguistically diverse grade-school math problems. These human-written problems need solutions that involve multi-step reasoning and a series of basic arithmetic operations. Second, MATH500 Hendrycks et al. (2021) contains 500 challenging problems sourced from high school math competitions with five distinct levels based on the Art of Problem Solving (AoPS) framework, ranging from level 1 to level 5. Third, AIME AIME is a math problem dataset collected from the American Invitational Mathematics Examination (AIME) competition from 1983 to 2024, designed to challenge the most exceptional high school math students in the United States. These problems cover various fields, such as algebra, geometry, and number theory.
Models. We evaluate our algorithm using three popular models: DeepScaleR-1.5B-Preview [^8] https://pretty-radio-b75.notion.site/DeepScaleR-Surpassing-O1-Preview-with-a-1-5B-Model-by-Scaling-RL-19681902c1468005bed8ca303013a4e2, DeepSeek-R1-Distill-Llama-8B DeepSeek-AI (2025), and Qwen1.5-MoE-A2.7B-Chat [^9] https://qwenlm.github.io/blog/qwen-moe/. These models span diverse architectures (dense vs. MoE), training recipes, and parameter scales. In this paper, we use DSR to denote DeepScaleR-1.5B-Preview, DSL to denote DeepSeek-R1-Distill-Llama-8B, and Qwe to denote Qwen1.5-MoE-A2.7B-Chat.
Metrics. We evaluate efficiency and accuracy using two metrics. Token savings is defined as the ratio of the number of tokens generated without Guardian to that generated after applying Guardian. Accuracy Wang et al. (2024) measures the mathematical equivalence between an LLM’s output and the ground-truth answer. For each test case, it is either correct or incorrect, and the overall accuracy is reported as the percentage of correctly solved problems across the entire dataset. When reporting accuracy changes, we use the absolute difference in accuracy, expressed in percentage points.
@@ 6.2 Evaluation Results
Cost of LZ77. We generate random strings of up to 128k tokens and measure the cost of LZ77 compression, which is approximately 34 ms. Because random strings exhibit minimal repetition and are more expensive to compress than natural language text, and because 128k tokens far exceed the sequence lengths used in our experiments, this measurement represents an upper bound. The resulting cost is comparable to the latency of decoding a single token Zheng et al. (2024). With $f=250$, compression is invoked once every 250 decoding steps, making the overall overhead negligible.
@@ 7.1 Sparse-Attention Algorithms
Sparse-attention algorithms exploit the fact that only a small subset of tokens receives significant attention scores, and we categorize them along two dimensions. (1) Inference stage. Some algorithms primarily accelerate the prefill stage, such as MInference Jiang et al. (2024), FlexPrefill Lai et al. (2025), and SpargeAttention Zhang et al. (2025), while others focus on the decode stage, including H2Zhang et al. (2023), StreamingLLM Xiao et al. (2024b), Quest Tang et al. (2024), InfLLM Xiao et al. (2024a), DuoAttention Xiao et al. (2025), and RaaS Hu et al. (2024). (2) Training dependency. Training-aware algorithms incorporate sparsity into the model architecture and are trained jointly with the model, such as DeepSeek NSA Yuan et al. (2025) and DSA Liu et al. (2024). In contrast, training-free algorithms operate as plug-ins and can be readily applied to pretrained models; examples include H2O, StreamingLLM, Quest, InfLLM, DuoAttention, and RaaS.
@@ 7.2 Chain-of-Thoughts Compression
Prior work on CoT compression, often referred to as long-to-short reasoning, can be broadly categorized into two classes. Training-aware approaches integrate compression into the training process, including ThinkPrune Hou et al. (2025), DLER Liu et al. (2025), and O1-Pruner Luo et al. (2025). In contrast, post-training approaches operate on pretrained models, such as Answer5 Liu and Wang (2025), HALT-COT Laaouach (2025), and UnCertT Zhu et al. (2025).
@@ 8 Conclusion
Evaluation on a limited set of datasets and models. Our evaluation covers only three models and three datasets. As such, the results may not generalize beyond these specific configurations. Although models with longer context lengths (e.g., Qwen2.5-Max, DeepSeek-r1) and datasets such as GPQA Diamond and Codeforces exist, exhaustive evaluation across all combinations is computationally prohibitive Hu et al. (2023). As reported in prior work Zhong et al. (2024), decoding a single token can take approximately 30 ms; thus, processing 16k tokens on an A100-80GB GPU requires around 8 minutes. Running 200 test cases would take over a day on a single GPU, making large-scale evaluation infeasible with limited resources. Despite these constraints, we select datasets spanning three levels of difficulty and models covering diverse architectures (dense vs. MoE), training recipes, and parameter scales. We therefore believe that the Lil problem is not an artifact of specific configurations but is universal across datasets and models.
Evaluation on a limited set of sparse-attention algorithms. Our evaluation covers only four sparse-attention algorithms, and thus the results may not directly generalize beyond this set. Nevertheless, we contend that the Lil problem is inherent to sparse-attention algorithms, as all assume sparsity patterns—i.e., that a few tokens are important—and may lose information (Section 4). We select diverse algorithms, including the pioneering H2O, the widely-used streamingLLM (recently applied in GPT-OSS OpenAI (2025)), the state-of-the-art Quest, and its counterpart infLLM. Other algorithms, such as clusterKV and PQCache, differ only in how K vectors are grouped and do not alter the use of sparsity. Hence, we believe the Lil problem is not specific to configurations but universal across sparse-attention algorithms.